OpenStudy (anonymous):

A ball is kicked into the air at time t=0. While the ball is in the air, its height above the ground after t seconds is -16t^2+40t feet. How many seconds after the ball is kicked will it hit the ground?

OpenStudy (kropot72):

Factorise the expression as follows: -16t^2 + 40t = 8t(-2t + 5) = 0 Can you see for what values of t the equation is true?

OpenStudy (anonymous):

Thank you for responding. What exactly do you mean by that?

OpenStudy (kropot72):

What values of t will make 8t(-2t + 5) equal to zero. One value of t that solves this equation is obviously t = 0. Can you work out what the other value is?

OpenStudy (kropot72):

The other value for t is found by solving the following: -2t + 5 = 0 Can you do this?

OpenStudy (anonymous):

Yes, I got it. Thank you very much!

OpenStudy (anonymous):

t=5/2

OpenStudy (kropot72):

-2t + 5 = 0 Subtract 5 from each side: -2t + 5 - 5 = 0 - 5 -2t = -5 t = -5/-2 = 2.5 seconds The ball will hit the ground 2.5 seconds after it is kicked into the air.

OpenStudy (kropot72):

Well done katlvb21 :)